OpenStudy (anonymous):

Find the lim as x-> 0 for 3x^4 - 6x^3/ 4x^3 + 2x^2

OpenStudy (anonymous):

Please help me, I know you can't plug 0 in so I'm confused on what to do now

OpenStudy (psymon):

\[\frac{ 3x^{4}-6x^{3} }{ 4x^{3}+2x^{2} }=\frac{ 3x^{3}(x-2) }{ 2x^2(2x+1) }=\frac{ 3x(x-2) }{ 2(2x+1) }\] Often times you need to try factoring. Take what I left ya and see what you can do :3

OpenStudy (anonymous):

But you still cant plug 0 into it yet, right? Becuase you'll get 0/2 which is undefined

OpenStudy (anonymous):

oh wait is the limit 0?